# Enhancing Object Detection Accuracy in Underwater Sonar Images through Deep Learning-based Denoising

### v1

@@ V Experiments
In this section, we conduct experiments based on the method proposed in Section 3 to evaluate the following two questions.
1.Whether applying denoising algorithms originally designed for optical images to sonar image datasets can improve object detection accuracy.
2.Whether treating the denoised dataset as multi-frame images and applying a multi-frame denoising algorithm can further improve object detection accuracy.
V-A Experimental Settings
To fairly evaluate the proposed questions, we utilized open-source underwater sonar image datasets, including FDD[39], UATD[38], SCTD[40], MarineDebris[41], and 2020Contest[42]. In SCTD [40], there are three types of targets: 461 shipwrecks , 90 aircraft, and 45 human figures. The contours and details of SCTD images are clear; however, interferences such as fragmentation, fractures, and geometric distortions can introduce additional challenges for detection. The UATD [38] dataset was collected using the Tritech Gemini 1200ik sonar from lakes and shallow waters. It contains over 9,000 multi-beam forward-looking sonar (MFLS) images. The 2020Contest [42] dataset includes both forward-looking and side-scan sonar images. The training set contains 2,288 images, with 1,216 side-scan sonar images, 800 forward-looking sonar images, and 272 negative samples with no targets. We selected the forward-looking sonar images for the experiment, as they provide a real-time, directional view of the forward environment, making them suitable for rapid detection and obstacle avoidance tasks, while side-scan sonar images are more suited for large-area, planar surveys, with lower accuracy and real-time performance. The Marine Debris dataset [41] was collected in the Ocean Systems Lab Water Tank at Heriot-Watt University, using an ARIS Explorer 3000 Forward-Looking Sonar operating at a 3.0 MHz frequency. The dataset includes various categories of marine debris, such as bottle, can, chain, drink carton, hook, propeller, shampoo bottle, standing bottle, tire, and valve. Each image is annotated with bounding boxes, and the annotations are stored in a corresponding JSON file. The Forward Looking Sonar Detection Dataset [39] is collected using the Oculus M1200d sonar, and it includes three categories: boat, plane, and victim, with a total of 3,752 images.
We chose three denoising methods [2][5][6] to comprehensively validate whether denoising techniques can improve object detection accuracy in underwater sonar images. Neighbor2Neighbor [2] provides a self-supervised approach to training an image denoising model using only noisy images, avoiding the need for large-scale clean-noisy image pairs. This method focuses on generating effective training pairs and reducing noise while maintaining the similarity of pixel appearances, making it suitable for underwater sonar images with varying noise distributions. Blind2Unblind [5] addresses the issue of information loss in self-supervised denoising by using a global-aware mask mapper and a re-visible loss function, which allows the network to capture important details in noisy images. This method performs well on both synthetic and real-world datasets, showing superior performance in denoising underwater sonar images. DSPNet [6] addresses the problem of speckle noise in sonar imaging, which significantly affects object detection. By using a feature pyramid network (FPN) and atrous spatial pyramid pooling (ASPP), DSPNet effectively reduces speckle noise while preserving key image details. The multi-scale mixed loss function further enhances its performance, making it suitable for underwater sonar images.
In the multi-frame denoising approach, we chose the multi-frame denoising algorithm OPD [65] because it provides a unified, unsupervised method that effectively leverages multiple noisy frames to enhance image quality. OPD learns through mutual supervision among frames, capturing inter-frame correlations and reducing noise without relying on clean labels. This strategy ensures adaptive information preservation, allowing critical details in different frames to be retained while enhancing robustness against noise.
We conducted object detection experiments using MMDetection, an open-source object detection toolbox developed by OpenMMLab, implemented based on PyTorch. This library provides a highly convenient platform for researchers and engineers to train and infer models.
To comprehensively evaluate the impact of denoising algorithms on object detection performance in underwater sonar images, we carefully selected four representative object detection algorithms: SSD-MobileNet v2, Faster R-CNN, SSD300, and YOLOX. These algorithms cover a diverse range of detection paradigms, enabling a robust assessment of how denoising influences detection accuracy, robustness, and computational efficiency.
First, SSD-MobileNet v2 is included due to its lightweight structure and efficiency, making it highly suitable for real-time applications on resource-constrained devices. Its use of MobileNet v2 as a backbone provides a balanced trade-off between detection accuracy and computational cost. This allows us to assess how denoising improves the performance of lightweight models designed for practical deployment in mobile or embedded systems.
Second, Faster R-CNN is chosen for its role as a representative of two-stage detection frameworks. Its region proposal network (RPN) enables precise detection and localization of objects, especially in scenarios with complex backgrounds or overlapping targets, such as in sonar images. By including Faster R-CNN, we aim to evaluate how denoising affects high-accuracy models that prioritize precision over speed.
Third, SSD300, as a single-stage detection algorithm, bridges the gap between speed and accuracy. With its ability to handle multi-scale feature maps for detecting objects of varying sizes, SSD300 is particularly relevant for sonar images, where targets may vary in size and clarity due to environmental noise. This inclusion helps us analyze the impact of denoising on models that balance real-time performance and moderate precision.
Finally, YOLOX represents the latest advancements in one-stage object detection, offering improved accuracy and speed compared to traditional YOLO models. Its advanced feature extraction and anchor-free design make it highly effective for detecting objects in challenging environments. Using YOLOX enables us to investigate the influence of denoising on state-of-the-art models that excel in both speed and detection performance.
By evaluating these four algorithms, we can comprehensively study the interplay between denoising and object detection across various detection paradigms. This selection ensures that our analysis captures the effects of denoising on models with different computational demands, detection strategies, and target applications, thereby providing valuable insights for improving object detection in underwater sonar imagery.
The experiments were conducted on a PC running Ubuntu 22.04.5 LTS operating system. The PC is equipped with an Intel(R) Core(TM) i9-14900K processor with 32 core. It has 128 GB of RAM and two NVIDIA GeForce RTX 3090 GPUs, each with 24 GB of memory, totaling 48 GB. The GPU driver version is 535.183.01, and the CUDA version used is 12.2.
Our experiments can be broadly divided into three parts. The first part involves performing object detection on the original underwater sonar datasets. The second part focuses on object detection on the denoised underwater sonar datasets. The third part conducts object detection on the underwater sonar datasets after multi-frame denoising.
V-B object detection on the original underwater sonar datasets
The detection results on the five datasets, FDD, 2020contest, MarineDebris, SCTD, and UATD, using four detection algorithms with the MMDetection toolkit are shown in Figures 3–7.
The analysis of the experimental results on five underwater sonar datasets reveals significant performance variations across different object detection algorithms, highlighting the critical impact of noise characteristics and target features in sonar images. For the UATD and FDD datasets, all algorithms exhibit high detection performance, with YOLOX and Faster R-CNN achieving remarkable mAP@0.5:0.95 scores of 0.545 and 0.845, respectively. These results demonstrate the potential of deep learning models in scenarios with relatively clear target features. Specifically, on the FDD dataset, YOLOX achieved the highest performance at high IoU thresholds (mAP@0.75 of 0.974) while maintaining a high AR, indicating a balanced trade-off between precision and recall. This superior performance is likely attributed to the distinctive target features and lower noise levels in the UATD and FDD datasets, which facilitate the models’ ability to extract and utilize critical information effectively.
In contrast, the performance of all algorithms declines significantly on the Marine and 2020contest datasets, with much lower mAP@0.5:0.95 scores compared to UATD and FDD. This decline can be attributed to higher noise levels, more complex backgrounds, or smaller target sizes, which collectively increase the difficulty of detection. On the MarineDebris dataset, YOLOX achieves the best performance in mAP@0.5 (0.96), but it lags behind Faster R-CNN in mAP@0.75 and AR, suggesting that while YOLOX excels at detecting targets under low IoU thresholds, it may struggle with precise boundary detection under stricter IoU conditions. On the 2020contest dataset, Faster R-CNN and YOLOX achieve comparable mAP@0.5:0.95 scores of 0.271 and 0.278, respectively. However, YOLOX’s advantage in AR indicates stronger recall capabilities, whereas Faster R-CNN appears more adept at detecting high-confidence targets in complex noise conditions.
On the SCTD dataset, the overall performance of all algorithms is relatively low, indicating the presence of extreme noise levels or blurred target features, which further complicates detection. On this dataset, SSD300 achieves the highest mAP@0.75 score (0.61), demonstrating its capability for detection under moderate IoU thresholds. Although YOLOX maintains some advantages in mAP@0.5, its mAP@0.5:0.95 performance is slightly inferior, reflecting potential limitations in its generalization to multi-scale targets in high-noise, low-contrast scenarios. These findings highlight the distinctive strengths of different algorithms when dealing with complex sonar data, emphasizing the importance of selecting appropriate algorithms for specific datasets. Furthermore, the results underline the critical challenge of balancing noise reduction and target feature preservation in underwater sonar object detection.
V-C object detection on the denoised underwater sonar datasets
We reviewed the state-of-the-art deep learning denoising algorithms and ultimately selected three that met our experimental requirements, enabling us to carry out a comprehensive experiment. The three deep learning-based denoising algorithms we selected each have unique strengths that meet our diverse experimental requirements for denoising performance. To begin with, Neighbor2Neighbor [2] achieves effective noise reduction by training on pairs of neighboring pixels within the same noisy image. This approach not only reduces noise significantly but also enhances model performance through a regularizer, thereby minimizing dependency on noise distribution assumptions. This enables it to achieve strong denoising without the need for clean images, making it robust for real-world noisy data. Additionally, Blind2Unblind [5] introduces an innovative transition from blind to non-blind denoising. By employing a global perception mask mapper and a re-visible loss, it improves denoising performance, allowing for precise noise removal while preserving critical image details. This balance between noise reduction and detail preservation makes it highly effective in scenarios requiring high fidelity. Finally, DSPNet [6] is specifically designed for the reduction of specks of noise. It applies a logarithmic transformation to convert multiplicative noise into an additive form, while leveraging feature pyramid networks and atrous spatial pyramid pooling to improve image quality and retain essential details. Each of these algorithms offers distinct strengths, providing a comprehensive basis for our experiments and comparative analysis of denoising performance.
The Blind2Unblind algorithm can be divided into four parameters: (1) Gaussian noise with , (2) Gaussian noise with , (3) Poisson noise with , and (4) Poisson noise with . A denoising model can be trained for each type of noise and its corresponding parameters. Similarly, the Neighbor2Neighbor algorithm can also be divided into four parameters: (1) Gaussian noise with , (2) Gaussian noise with , (3) Poisson noise with , and (4) Poisson noise with . In addition to these, the denoising model trained by DSPNet is also included. In total, there are nine denoising models. In the inference stage, the dataset is input into the trained models to obtain the denoised underwater sonar image datasets.
Tables 2–29 present the detection results of five datasets evaluated across four object detection algorithms, including mAP (), mAP_50 (), mAP_75 (), mAP_s, mAP_m, mAP_l, AR, AR_s, and AR_l, where s, m, and l represent small targets (objects with areas smaller than ), medium targets (areas between and ), and large targets (areas larger than ), respectively. The term ”noisy” denotes detections performed on the original dataset without denoising.

### v2

@@ 5 Experiments
In this section, we conduct experiments to address the following key research questions:
1.
Can traditional optical denoising methods generalize to sonar data without modification?
2.
Which denoising model families are most robust to sonar noise patterns?
3.
Can the accuracy of target detection be further improved by the proposed multi-source denoising fusion algorithm?
5.1 Datasets
To evaluate our denoising and detection framework, we selected five publicly available sonar datasets covering diverse underwater scenarios, target types, and noise conditions. Below, we outline their key characteristics and relevance to our research objectives.
1.
SCTD (Sonar Composite Target Dataset) [17]: Contains 596 annotated targets (461 shipwrecks, 90 aircraft, 45 human figures) with high-resolution contours, facilitating quantitative analysis of denoising effects on structured objects. The dataset’s fragmented and geometrically distorted images simulate real-world sonar artifacts, testing algorithm robustness.
2.
UATD (Underwater Acoustic Target Dataset) [16]: Comprises >9,000 multi-beam forward-looking sonar (MFLS) images captured in shallow waters using a Tritech Gemini 1200ik sonar. Its large-scale, real-world data is essential for assessing generalizability across varying water conditions and target occlusions.
3.
USD (URPC_sonarimage_dataset) [19]: Includes 800 forward-looking and 1,216 side-scan sonar images. We focus on the forward-looking subset due to its real-time directional perspective, which aligns with dynamic obstacle avoidance applications.
4.
MarineDebris [18]: Collected in a controlled water tank using an ARIS Explorer 3000 sonar (3.0 MHz), this dataset features 10 fine-grained debris categories (e.g., bottles, tires, propellers) with precise bounding-box annotations (JSON format). It enables rigorous evaluation of denoising’s impact on small-object detection in cluttered environments.
5.
FDD (Forward-Looking Sonar Detection Dataset) [15]: Contains 3,752 images of boats, planes, and victims, acquired with an Oculus M1200d sonar. Its balanced class distribution and standardized annotations allow direct comparison of detection accuracy improvements after denoising.
5.2 Infrastructure Configuration
All experiments were implemented using MMDetection (PyTorch-based) and conducted on a workstation with:
1.
CPU: Intel Core i9-14900K (32 cores)
2.
GPU: 2× NVIDIA RTX 3090 (24 GB each, CUDA 12.2)
3.
RAM: 128 GB
4.
OS: Ubuntu 22.04.5 LTS
5.3 Detection on Raw Sonar Images
We first established detection performance baselines by evaluating four representative architectures (SSDMobileNetV2, Faster R-CNN, SSD300, and YOLOX) on unprocessed sonar images using the MMDetection framework. This controlled experiment serves to quantify the intrinsic detection capability of each model in challenging underwater noise conditions, with all models trained and tested using identical dataset partitions to ensure methodological consistency.
The training procedure followed each detector’s standard recommended configuration, including optimizer selection and epoch count. For comprehensive performance evaluation, we employed multiple established metrics: mAP@[0.5:0.95] as the primary measure of overall detection quality, supplemented by mAP@0.5 and mAP@0.75 for IoU-specific accuracy assessment, along with Average Recall (AR) to evaluate detection robustness. These baseline measurements precisely characterize each detector’s performance limitations when processing noisy sonar imagery directly.
Fig. 4 presents the baseline object detection performance evaluation conducted on raw sonar images across five underwater datasets using four representative detection algorithms. The results reveal significant variations in model performance depending on dataset characteristics. All models achieved their highest mAP@0.5:0.95 scores (ranging 0.6-0.8) on the FDD dataset, which benefits from balanced class distribution and standardized annotations. In contrast, performance degraded substantially on USD and SCTD. The most significant performance variations occurred in the SCTD and UATD datasets, exhibiting the greatest model instability across detection methods.
5.4 Benchmark Evaluation of Object Detection in Sonar Images Using Different Denoisers
This experimental study evaluates object detection performance on sonar images processed through nine denoising variants derived from three fundamental methods. The comprehensive results, presented in Appendix Tables 3– 38, systematically compare detection accuracy using four algorithms (SSDMobileNetV2, Faster R-CNN, SSD300, and YOLOX) across five sonar datasets, where the “Noisy” condition represents baseline performance on unprocessed images. The evaluation employs nine rigorous metrics: mean Average Precision (mAP) at different Intersection-over-Union thresholds (mAP@, mAP@, and mAP@)), along with scale-specific variants (mAPs, mAPm, mAPl, AR, ARs, and ARl). Here, the subscripts denote target size categories: small ( pixels), medium ( pixels), and large ( pixels) targets, enabling detailed analysis of detection performance across varying target scales in sonar imagery.